OpenStudy (anonymous):

Consider this equation: 4x – 5 + 6x = 6 Which is an equivalent equation after combining like terms? A. -1 + 6x = 6 B. 10x - 5 = 6 C. 4x + 1 = 6 D. 4x - 5 = 0

OpenStudy (kva1992):

10x-5=6

OpenStudy (anonymous):

B. 10x-5=6, because combining 6x+4 will give you 10x-5=6